OpenStudy (anonymous):

Nancy left a bin outside in her garden to collect rainwater. She notices that 1/2 gallon of water fills 2/3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words. I don't want answer i want help.

OpenStudy (anonymous):

ok i will walk you through it

OpenStudy (anonymous):

r u ready

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Sorry i have a slow connection right now..

OpenStudy (anonymous):

But im ready.

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

if 1/2 filled 2/3 the we can write an equation stating the relationship betwween 1/2 and 2/3

OpenStudy (anonymous):

1/2 = 2/3 right. So how do we determine this? i might be over thinking.

OpenStudy (anonymous):

so 1/2 = 2/3 of bin right

OpenStudy (anonymous):

ok therefore 1 water is now 4/3 right

OpenStudy (anonymous):

of bin

OpenStudy (anonymous):

Yes? im sorry these things confuse me. Fractions suck.

OpenStudy (anonymous):

but the question says we want one of bin so what do you do you cross multiply

OpenStudy (anonymous):

3/4 water = 1 of bin do you get it or i should come again

OpenStudy (anonymous):

its a essay not a multiple choice so i dont know how to say it is the problem...

OpenStudy (anonymous):

ok so i should put it in a better format

OpenStudy (anonymous):

let the water be x and bin =y

OpenStudy (anonymous):

then 1/2 x=2/3y therefor for us to gey 1 y (filled up,) we cross multiply

OpenStudy (anonymous):

3*1/2*2 = 3/4

OpenStudy (anonymous):

im done do you get it now

OpenStudy (anonymous):

Think so. Thanks alot!

OpenStudy (anonymous):

you welcome